Is the function f defined by f(x) = tan(2x) is a periodic function with period π

1
Expert's answer
2021-11-01T17:40:10-0400

Taking into account that the smallest period of the function "\tan(x)" is "\pi," we conclude that the smallest period of the function "f(x) = \tan(2x)" is "\frac{\pi}2." Since "\pi=2\cdot\frac{\pi}2," we conclude that "\pi" is also a period of "f," but it is not the smallest period.